OpenStudy (anonymous):

Use substitution to solve this linear system : x+3y=7 5x-4y=-10

OpenStudy (anonymous):

(1) x+3y=7 (2) 5x-4y= -10 from (1) we have: x = 7 - 3y (1.1) Plug (1.1) into (2) and we have: 5(7-3y)-4y= -10 35-15y-4y = -10 35 - 19y = -10 - 19y = -10 - 35 - 19y = -45 y = (-45) / (-19) = 45/19 use the result to find x with (1.1) x = 7 - 3(45/19) x = 7 - 135/19 x = (133-135) / 19 = -2 / 19 is that right?